OpenStudy (anonymous):

What is the slope of the line that passes through the given points? (2, 12) and (6, 11) A. -1/4 B. 1/4 C.4 D.-4

OpenStudy (anonymous):

y=-0.25x+12.5

OpenStudy (anonymous):

to find the slope use this formula\[\frac{ (y _{2}-y _{1}) }{ (x _{2}-x _{1}) }\]

OpenStudy (anonymous):

it doesnt matter what you choose as coordinate 1 or 2, but stay consistent

OpenStudy (anonymous):

first find the gradient, and use that to form the equation using, y-y1=m(x-x1)

OpenStudy (anonymous):

Hint: Find slope \[slope = \frac{ y2-y1 }{ x2-x1 }\] Use one of the points and put into POINT-SLOPE form. y - y1 = m(x - x1) m = slope Convert to SLOPE-INTERCEPT form. y = mx + b